OpenStudy (anonymous):

The area of a rectangle is given by the expression x2 + 5x + 4. If the length of one side is given by x + 2, what is the length of the other side?

OpenStudy (jdoe0001):

Area = Length * Width if the Length = x+2 and we know the Area, so \(\bf Area= x^2 + 5x + 4\qquad Length = x+2\\ \quad \\ Area = Length \cdot Width\implies x^2 + 5x + 4 = (x+2)\cdot Width\\ \quad \\ \cfrac{x^2 + 5x + 4}{x+2}=Width\)